find solution to ivp

xy' + 3y = x(sqrt(y)), y(1) = 0, bernoulli equation
Answer #1


1 To solve my +3y = x1, y(1) = 0 Divide the above equation with x, we get 3 y+-y= y2... (1) Clearly (1) is the Bernoullis

The general solution of (3) is 24(x) = [ 4(x)Q(x) dx+c (2)=1*()to 3 x²dxto 3 1 x2 23 2 +c + = x3+ S = -x +c 5 = zx x +c 5 13